# Experience-Driven Exploration for Efficient API-Free AI Agents

### v2

@@ 3.3 Reasoning and Decision-Making Module
Figure 3: Game progression measured by Civilization V’s technology tree. Our framework outperforms all baselines, including those with task-related priors.
4 Experiments
Experimental Setup. We evaluate KG-Agent in two challenging games: Slay the Spire and Civilization V. Both games require not only low-level execution accuracy but also high-level long-horizon planning, making them ideal testbeds for assessing strategic reasoning and adaptation. Their turn-based nature also allows seamless integration with LLM-based reasoning under current latency constraints. Due to the limitations of existing frameworks, which often rely on predefined APIs, task-specific prompts, or are ineffective even with prior knowledge (e.g., subgoals or game rules), we compare against several representative baselines: GPT-4o, Claude 3.7, the open-source model UI-TARS-1.5 Qin et al. (2025), and the Bottom-up Agent Du et al. (2025). Specifically, we use GPT-4o for semantic understanding and reasoning, and CLIP ViT-B/32 as the visual encoder. All hyperparameters remain consistent across experiments. For the SA-KG, the state merging threshold $\theta_{\text{merge}}$ is set to 0.95, and the similarity edge threshold $\theta_{\text{simi}}$ to 0.88. For skill edge weight computation (Eq. 1), we use a balancing factor $\alpha=0.7$ and fitness sensitivity constant $C_{0}=5.0$. The UCT-based skill selection (Eq. 6) uses an exploration constant $C_{1}=5.0$. All environments share the same agent implementation, demonstrating the generality of our API-free framework. Each agent runs for three episodes per environment, ending upon game completion, failure, or after $500$ steps.
Table 2: Evolution of skill library and SA-KG, as well as game performance over training rounds in Civilization V. Each round builds on previous memory with 100 steps.
Round Skill Library Information SA-KG Information Civilization V
[…]
Round 0 76 77 1 26 37 226 65 10 0.89 3.0
Round 1 106 33 3 50 84 782 98 12 0.92 2.8
Round 2 111 7 2 53 90 856 113 13 0.93 2.7
Round 3 114 4 1 55 90 838 115 13 0.94 2.7
@@ 4 Experiments
((c))
((d))
Figure 4: Visual evolution of the SA-KG from Round 0 to Round 3. Skill associations (red) and similarity links (blue) illustrate structural refinement and relational reasoning across rounds.
Table 3: Ablation study on the contribution of the experience neighborhood and the SAKG-driven reward, with each run constrained to 100 steps for clarity. The “w/o similarity edge” means that each state forms an isolated node, with no merging or similarity-based connections permitted.
[…]
Size Size Edges Edges (Turns) Research. Respons. Rate Costs
Full KG-Agent Model 76 26 37 226 65 10 0.89 3.0
   w/o similarity edge 30 672 62 0 15 1 0.64 3.5
   w/o $\mathcal{R}_{novel}$ 48 11 19 74 23 2 0.81 3.1
   w/o $\mathcal{R}_{state}$ 57 20 33 110 39 4 0.87 3.2
   w/o $\mathcal{R}_{novel}$ & $\mathcal{R}_{state}$ 64 18 39 96 35 3 0.79 3.3
((a))
((b))
@@
Figure 5: Skill Invocation and critical skills. (a) Our framework executes high-reward skills even with minimal visual change between states. (b) The top 10 skill edges reflect a consistent focus on critical long-term actions, aligning with the game’s core progression mechanics.
4.1 Comparative Results
As shown in Table 1 and Figure 3, the KG-Agent demonstrates superior performance across two distinct open-ended games compared to all baseline agents, regardless of whether they utilized prior assistance. Notably, our agent attains a 99% and 94% execution-responsive rate, indicating that our agent achieves significantly higher in-game progression and scores while maintaining a similarly high execution-responsive rate. Critically, our approach accomplishes this with reduced computational overhead, as evidenced by lower token costs per 100 steps compared to BottomUp agent. These results indicate that our agent not only inherits the BottomUp agent’s reliability in skill execution but also enhances strategic decision-making, enabling more effective long-term planning and adaptation in open-ended gameplay. The combination of superior performance and improved efficiency underscores the effectiveness of our approach in mastering complex game environments.
4.2 Ablation Studies
Skill Evolution over Rounds. Table 2 and Figure 4 illustrate the evolution of the skill library and SA-KG over four training rounds in Civilization V. The skill library grows dynamically through iterative augmentation and pruning, reflecting continuous refinement. Concurrently, the SA-KG expands in structural complexity, with similarity edges increasing notably, highlighting enhanced relational reasoning. These improvements correlate with steady gains in game progression and execution responsiveness, while token costs decrease, indicating improved reasoning efficiency. The convergence of metrics by Round 3 confirms that repeated skill reuse and structural consolidation facilitate sustained agent adaptation and strategic exploration in open-ended environments.
Effect of Core Modules. As shown in Table 3, the removal of similarity edges severely impairs long-term progression, while disabling the reward mechanism reduces strategic coherence, collectively underscoring their necessity for sustaining effective open-ended gameplay. The efficacy of these components is further validated by the SA-KG operational behavior. Specifically, our framework demonstrates its ability to prioritize strategic value over perceptual salience by invoking high-reward skills such as “Advance Tur” and “Move Worker” even when visual changes between states are minimal, as evidenced in Figure 5(a). This validates the reward mechanism’s capacity to identify critical actions beyond superficial cues. Meanwhile, the key skills identified by the SAKG, quantified in Figure 5(b), reveal a consistent strategic focus. Skills with the highest edge weights, e.g., “Adopt Policy Action” (0.6776) and those related to unit and worker management, are precisely those that drive long-term progression, reflecting the critical role of the experience neighborhood in maintaining behavioral consistency and a coherent long-term strategy.
5 conclusion
@@ 4.2 Ablation Studies
In this paper, we proposed KG-Agent to organizes pixel-level GUI interactions into a knowledge graph of states and actions, enabling agents to overcome short-sightedness in API-free, open-ended environments. By linking functionally similar yet visually distinct states, the agent generalizes past experiences into coherent long-term strategies. Experiments in open-ended environments demonstrate that KG-Agent achieves significant gains in exploration efficiency and strategic depth over state-of-the-art baselines. These results highlight the potential of structuring experience into a graph-based memory, advancing API-free agents toward scalable and general-purpose autonomy.
@@
Limitations and future work. A key limitation of KG-Agent is that its knowledge remains tied to specific environments, lacking the full abstraction and transferability that characterize human reasoning. While the framework structures experience into a graph-based memory, its ability to generalize knowledge across unseen tasks and diverse contexts is still constrained. Future work will focus on advancing this direction: enabling higher-level knowledge induction and reasoning that can span heterogeneous environments and ultimately testing such capabilities in more complex, real-world scenarios. This progression would move API-free agents closer to human-like adaptability and robust general intelligence.
Appendix A Appendix
A.1 More Related Works
LLM-Based Multi-Agent Systems. The integration of LLMs Achiam et al. (2023); Liu et al. (2024a); Bai et al. (2023) into multi-agent systems has enabled a new architectural paradigm in which agents actively reason, communicate, and interact with external tools, environments, and users through structured interfaces Sun et al. (2024); He et al. (2025); Li et al. (2024). By leveraging APIs and GUI-based interactions, these systems elevate LLMs from passive text generators to dynamic participants capable of orchestrating complex, multi-step tasks Zhang et al. (2024a); Tzachristas (2024); Liu et al. (2024b). In practice, modern LLM-based agents function as autonomous or semi-autonomous entities that combine natural language understanding with tool-augmented reasoning Zhao et al. (2025); Wang et al. (2025). Moreover, GUIs further extend an agent’s interactive capacity Tang et al. (2025a). In simulated or game environments, GUI perception modules can translate visual elements, e.g., buttons, maps, menus, into semantic representations that an LLM can reason about Hu et al. (2025). This approach enables LLM agents to operate across diverse digital environments, from software applications to strategy games, without requiring task-specific policy training Tang et al. (2025b); Liu et al. (2025). Notably, frameworks such as AutoGen Wu et al. (2024), MetaAgent Li et al. (2023), and Generative Agents Park et al. (2023) formalize these patterns, offering abstractions for agent roles, tool integration, message routing, and environment interfacing. By unifying natural language communication with API-driven and GUI-mediated interaction, LLM-based multi-agent systems are advancing from conceptual models to deployable platforms for open-ended, human-aligned reasoning and decision-making. Recently, the application of LLM-based multi-agents to complex, multi-step tasks has yielded remarkable achievements across diverse domains Xi et al. (2025), e.g., web browsing Gu et al. (2024), software operation Jin et al. (2024), and robotics Firoozi et al. (2025).
API-Free LLM-Based AI Agents. While early LLM-based agents frequently depended on privileged access to internal game states via APIs, this API-centric paradigm faces critical constraints Wang et al. (2023a). Its reliance on engineered interfaces inherently limits generalization, particularly in closed-source commercial games and proprietary software where internal APIs are inaccessible or undocumented Tan et al. (2025). Moreover, such abstractions can distance the agent from the rich visual context and fine-grained control characteristic of genuine human-computer interaction. In response, a more general and challenging direction has emerged: API-free LLM-based agents. These systems operate purely from raw pixel input, interacting through universal human-style interfaces such as keyboard and mouse, without relying on internal state APIs or specialized function calls Tan et al. (2025); Du et al. (2025). This approach targets ultimate generality, enabling agents to operate across any visible software environment, much like a human user, without custom integration. Notable progress has been made under this paradigm. In desktop automation, multimodal agents such as UFO Zhang et al. (2024b) and CogAgent Hong et al. (2024) interpret GUI elements and execute multi-step workflows based on natural language instructions, marking a stride toward generalist assistive agents. In open-ended game environments, systems like CRADLE Tan et al. (2025) and the Bottom-Up Agent Du et al. (2025) have demonstrated that LLM-based agents can learn complex tasks from pixel-level input via autonomous trial-and-error, achieving meaningful in-game progress without game-specific adaptations. These advances underscore the potential of API-free agents to operate in broadly applicable and visually grounded settings, paving the way for more universal and flexible AI systems.
((a))
((b))
((c))
((d))
Figure 6: Structured Prompt Design for KG-Agent’s VLM Reasoning Stages. (a) Prompt for Skill Invocation enables action selection by analyzing the game state and action conditions. (b) Prompt for Skill Augmentation derives new actions from UI-level operations and screen changes, where $\mathbb{O}$ is the set of operations. (c) Prompt for Skill Refinement (Cluster and Merge) groups to functionally similar new actions. (d) Prompt for Skill Evaluation assesses action effectiveness based on expected versus actual game state changes.
A.2 More Methodological Details
Prompt Engineering for Agent Reasoning. Effective reasoning in visual-language models (VLMs) relies on meticulously designed prompts to guide complex visual understanding and decision-making processes Shtedritski et al. (2023); Gu et al. (2023). In KG-Agent, we employ a structured VLM-based reasoning cycle consisting of four key stages: skill invocation, augmentation, refinement, and evaluation. As illustrated in Figure 6, each stage is supported by a dedicated prompt template that collectively enables intelligent interaction in dynamic GUI environments without reliance on pre-defined APIs. Specifically, the $prompt^{invoke}$ in Figure 6(a) serves as the agent’s primary decision-making interface. It selects the most suitable executable action from procedural memory when symbolic reasoning over the SA-kG fails. The $prompt^{augment}$ in Figure 6(b) complements invocation by deriving new high-level actions from low-level UI operations ($\mathbb{O}$), thereby expanding the agent’s skill repertoire based on observed screen changes. The $prompt^{refine}$ in Figure 6(c) consists of two sub-prompts: one groups functionally similar new actions, and the other merges them with existing skill clusters to maintain a compact and non-redundant skill library. Finally, the $prompt^{evaluate}$ in Figure 6(d) assesses the effectiveness of executed skills by comparing expected outcomes against actual screen changes, providing essential feedback for progressive and semantic rewards as defined in Eq. 9. Together, these prompts form a closed-loop learning architecture: invocation drives action, augmentation acquires knowledge, refinement consolidates understanding, and evaluation validates progress. This integrated design ensures that KG-Agent can continuously perceive, reason, act, and adapt within complex, unseen GUI environments.
Pseudocode of The KG-Agent. All skill invocation, augmentation, refinement, and evaluation steps follow the algorithmic loop defined in Algorithm 1. A critical innovation of our approach is the identification of the Neighborhood of Experience $\mathcal{N}_{\mathcal{E}}(v_{i})$, which enables the agent to leverage past experiences in similar states. This neighborhood information facilitates the gathering of high-quality candidate skills $\mathbb{S}_{HQ}(v_{i})$ from the SA-KG, representing skills that have proven effective in comparable situations. The agent performs multiple attempts to execute them, with each attempt generating reward feedback that updates the skill edges in the SA-KG. The loop breaks early if a sufficiently high reward is achieved, ensuring efficient skill utilization. When no high-quality skills are available, the agent enters a trial-and-error phase. This integrated approach enables the KG-Agent to dynamically balance exploitation of proven strategies with exploration of novel solutions, while maintaining an evolving knowledge structure that captures both state relationships and skill effectiveness across the environment.
Algorithm 1 The Pseudocode of Our KG-Agent Skill Evolution
Input: Environment $(\mathcal{X},\mathcal{A},\mathcal{T},\mathcal{R})$, VLM $f_{VLM}$, Skill library $\mathbb{S}=\emptyset$, and SA-KG $\mathcal{G}(\mathcal{V,E})=\emptyset$.
1:for Each Agent do
2:  while Episode not terminated do
3:   Observe current screen $X_{i}\in\mathcal{X}$ and extract visual feature $x_{i}$;
4:   Generate or merge state node $v_{i}$ in $\mathcal{G}$, connect similarity edges $e^{sim}\in\mathcal{E}^{sim}$;
5:   Identify Neighborhood of Experience $\mathcal{N}_{\mathcal{E}}(v_{i})$ by Eq. 2;
6:   Gather high-quality candidate skills $\mathbb{S}_{HQ}(v_{i})$ from $\mathcal{G}$ by Eq. 3;
7:   if $\mathbb{S}_{HQ}(v_{i})\neq\emptyset$ then
8:    for $attempt=1$ to $max\_{attempts}$ do
9:     Sample high-quality candidate skill $\sigma_{k}$ from $\mathbb{S}_{HQ}$ by Eq. 4;
10:     Observe trajectory $\mathcal{T}$ and compute $\mathcal{R}_{total}(v_{i},v_{j})$ by Eq. 9;
11:     Generate skill edges $w^{\sigma_{k}}_{i,j}$ by Eq. 1;
12:     if $\mathcal{R}_{total}(v_{i},v_{j})>threshold$ then
13:      Break;
14:     end if
15:    end for
16:   else
17:    Sample candidate skill $\sigma_{l}$ from $\mathbb{S}_{C}(v_{i})$ by Eq. 7;
18:    if $\mathbb{S}_{C}(v_{i})=\emptyset$ then
19:     for $k=1$ to $k_{max}$ do
20:      Generate $a\leftarrow f_{VLM}(prompt^{augment},X_{i},\mathbb{O})$;
21:      if $a$ is recognizablerecognizable then
22:       Update Skill library $\mathbb{S}$;
23:       break;
24:      end if
25:     end for
26:    else
27:     Select the best skill cluster by $f_{VLM}(prompt^{invoke},X,\mathbb{S})$;
28:     Evaluate the skill cluster with MCTS;
29:     Execute the best skill $\sigma$ from MCTS;
30:    end if
31:    Observe trajectory $\mathcal{T}$ and compute $\mathcal{R}_{total}(v_{i},v_{j})$ by Eq. 9;
32:    if $\mathcal{R}_{total}(v_{i},v_{j})<threshold$ then
33:     Remove $\sigma$ from $\mathbb{S}$;
34:    end if
35:    if skill_count<thresholdskill\_count<threshold then
36:     Refine skill: $\sigma^{\prime}\leftarrow f_{VLM}(prompt^{refine},X,\sigma,\mathcal{T})$;
37:     Replace $\sigma$ with $\sigma^{\prime}$ if improvement observed;
38:    end if
39:   end if
40:  end while
41:end for
A.3 More Experimental Results
((a))
((b))
Figure 7: In-game settings and initial game states for Slay the Spire and Civilization V, maintaining consistent observation conditions and fair evaluation benchmarks.
Open-ended Environment. As shown in Figure 7, we evaluate the KG-Agent in two challenging games Slay the Spire (base difficulty: Ascension 0, character: Ironclad), which blends roguelike and deck-building mechanics, and Civilization V (Russia, Earth map, Standard size, Prince difficulty, Standard pace), a hallmark 4X (eXplore, eXpand, eXploit, eXterminate) strategy game.
Figure 8: In-game advancement measured by floors cleared and official run score in Slay the Spire.
((a))
((b))
Figure 9: The State-Action Knowledge-Graph (SA-KG) evolution from step 100 to 200 shows network growth ($22\rightarrow 28$ nodes, $190\rightarrow 312$ similarity edges, and $37\rightarrow 51$ skill edges), with skill (red) and similarity (blue) connections reflecting improved structural organization and relational reasoning capabilities.
More Experimental Results. Figure 8 illustrates the in-game advancement measured by floors cleared and official run score in Slay the Spire. Figure 9 illustrates the structural evolution of the SA-KG from step 100 to step 200, with the evaluation terminating at step 210 after the agent’s character is defeated by a monster. Over this interval, the knowledge graph undergoes substantial growth: the number of nodes increases from 22 to 28, similarity edges rise from 190 to 312, and skill edges grow from 37 to 51. This expansion reflects advancing relational reasoning capabilities—particularly the accelerated formation of similarity edges, which signal improved recognition of state equivalences and environmental patterns. Meanwhile, the steady increase in skill edges suggests effective knowledge transfer and progressive maturation of the graph, driven by denser connectivity and the emergence of cohesive clusters that support more robust decision-making.
Figure 10: Skill invocation with low visual change but high reward in Slay the Spire.
Figure 10 demonstrates our agent’s capacity in Slay the Spire to prioritize strategic value over perceptual salience, as evidenced by the selection of combat skills like “Ironclad Strike Sequence” (Total Reward: 2.114, Change_Ratio: 0.015) and “Attack and Select Strategy” (Total Reward: 2.086, Change_Ratio: 0.011) that yield consistent rewards despite minimal visual changes between states. This reflects the agent’s advanced reasoning capabilities in identifying strategically valuable actions—such as optimal card sequencing and targeted enemy selection—that contribute to incremental combat advantage and resource management, even when immediate visual feedback is subtle. The ability to transcend superficial interface cues and consistently execute high-value tactical decisions underscores the framework’s sophistication in modeling complex game dynamics, where subtle but strategically sound actions accumulate toward long-term success in challenging roguelike environments. In addition, Figure 11 also demonstrates KG-Agent’s effectiveness in overcoming the key limitations of API-free GUI agents by structuring pixel-level interactions into the SA-KG. The high-utility skill edges, e.g., combat sequences and post-battle navigation, show the agent’s ability to prioritize strategic actions beyond immediate rewards, overcoming myopic decision-making. Simultaneously, the frequent invocation of abstract skill clusters, particularly the dominant ”Select Ironclad Character” with 92 invocations, demonstrates successful generalization across visually distinct states, enabling efficient long-term planning by recalling proven strategies rather than relying on repetitive pixel-level exploration.
((a))
((b))
Figure 11: The structured skill acquisition of KG-Agent in Slay the Spire, demonstrating how the SA-KG framework overcomes pixel-level myopia. (a) High-utility skill transitions showing optimized combat sequences learned through the hybrid reward mechanism. (b) Frequently invoked skill clusters revealing hierarchical organization from basic operations to strategic planning, evidencing effective generalization across visually distinct GUI states.
A.4 Case Study
Here we present two representative case studies from the game Slay the Spire to illustrate further how KG-Agent structures pixel-based interactions into reusable skills and enables strategic generalization through its SA-KG.
Figure 12 presents a case study demonstrating successful skill reuse through the SA-KG of KG-Agent. When encountering a new state and corresponding node $v_{3}$, the proposed KG-Agent first retrieves its Neighborhood of Experience $\mathcal{N}\mathcal{E}(v_{3})$ and identifies high-quality candidate skills $\mathbb{S}{HQ}(v_{3})$. Skills are then sampled probabilistically according to their skill edge weights in the SA-KG. While the initially sampled skill “Use Defend Card in Battle” fails due to grounding issues, the subsequently sampled skill “Quick Attack Sequence” executes successfully through the simulated mouse operations “Clicked at (673, 926)” and “Clicked at (523, 951)”, achieving a total reward of 2.480 and transitioning to next state node $v_{6}$. Following successful execution, KG-Agent updates both the fitness value and edge weight of this skill in Memory, thereby reinforcing this validated strategy for future reuse. The case study in Figure 12 illustrates how KG-Agent addresses the key limitations of traditional LLM-based agents in API-free environments. First, by organizing pixel-level interactions into SA-KG, the agent overcomes inefficient exploration by generalizing across visually distinct but functionally similar states through Neighborhood of Experience, avoiding myopic decisions. Second, the edge-weight based sampling demonstrates strategic reuse of historical knowledge rather than relying on trial-and-error. Third, the successful execution and subsequent memory update illustrate our hybrid reward mechanism: the substantial environmental reward combined with fitness updates reinforces high-value pathways while maintaining exploration flexibility. This showcases KG-Agent’s ability to decouple strategic planning from pure discovery, effectively addressing both skill acquisition and long-horizon reasoning challenges in API-free environments.
Figure 12: Case study of successful skill invocation through state-conditioned sampling from the SA-KG. Given the current state, KG-Agent retrieves high-quality candidate skills from the neighborhood of experience $\mathcal{N}_{\mathcal{E}}$ and samples skills according to the weights of skill edges. While the first sampled skill “Use Defend Card in Battle” fails due to grounding issues, the second skill “Quick Attack Sequence” executes successfully. The agent subsequently updates the skill’s fitness and reinforces its edge weight in the Memory.
Figure 13 demonstrates KG-Agent’s two-stage skill invocation strategy, which prioritizes structured knowledge from the SA-KG before resorting to VLM-guided search. Initially facing state node $v_{6}$, the KG-Agent attempts SA-KG-based skill sampling but encounters five consecutive failures due to grounding issues or insufficient rewards. This triggers a fallback to VLM-guided skill retrieval from Procedural Memory. The process first queries the Action Clusters Table to obtain candidate skill clusters $\mathbb{S}_{C}(v_{6})=\{1,2,3,5,6,7,8,9,12\}$, from which the VLM identifies the most contextually relevant cluster based on visual observation. To navigate the stochastic environment, the KG-Agent employs a UCT-inspired strategy that balances exploitation of high-value skills $\{\sigma_{2},\sigma_{3},\sigma_{9},\sigma_{11}\}$ with exploration of less-tried options. The resulting utility scores are converted into a probability distribution via temperature-scaled softmax, enabling stochastic sampling. The first sampled skill “Attack and Select Strategy” (Probs.=9.99e01Probs.=9.99e-01) executes successfully through operations “Clicked at (687, 933)” and “Clicked at (1134, 692)”, yielding a 2.117 reward. This successful execution triggers updates to the skill’s fitness and edge weight in memory, demonstrating KG-Agent’s capacity for continuous knowledge refinement through both successful and failed experiences.
Figure 13: Case study of successful skill invocation by general VLM-guided trial and error. After five unsuccessful attempts with SA-KG-sampled skills, KG-Agent switches to trial-and-error mode. It first collects skill clusters from Procedural Memory using VLM-based selection, then applies MCTS with UCTS strategy to balance exploration and exploitation. The successfully executed skill “Attack and Select Strategy” is subsequently updated in memory with revised fitness and edge weight.
A.5 Large Language Models Usage Statement
In accordance with the conference policy on the use of LLMs, we disclose that an LLM was used solely as a general-purpose assistive tool for grammatical refinement and language polishing of the manuscript. The LLM played no role in research ideation, experimental design, data analysis, interpretation of results, or substantive writing. All scientific content, claims, and technical contributions were conceived, developed, and verified exclusively by the human authors. The authors take full responsibility for the accuracy, integrity, and originality of the work.

### v3

@@ IV-C Ablation Studies
Figure 4: The SAG evolution from step 100 to 200 shows network growth in Slay the Spire, with skill (red) and similarity (blue) connections reflecting improved structural organization and relational reasoning capabilities.
IV-D Model Analysis
Hyperparameter Sensitivity. For certain hyperparameters, e.g., the criterion for increasing skill length, the fitness sensitivity and greediness weighting constants, and the temperature for skill selection, we adopt values grounded in established practices from related work [13]. Furthermore, for the key hyperparameters directly tied to our core contributions, i.e., the number of skills sampled per execution attempt, the coefficient balancing state-value and novelty rewards, and the cosine-similarity thresholds for node merging and edge creation, we conduct a targeted sensitivity analysis in Slay the Spire. Specifically, for each of these key parameters, we evaluated two alternative values alongside the default setting, while keeping all other configurations identical. As shown in Table III, our selected hyperparameter configuration, sampling a maximum of 5 skills per execution attempt ($M=5$), applying a balanced reward coefficient ($\mathcal{R}_{3}:\mathcal{R}_{4}=1:1$), and setting cosine similarity thresholds for node merging and edge creation to $\theta_{merge}=0.95$ and $\theta_{simi}=0.88$, stands out as the optimal setup. It achieves superior in-game performance, attaining the highest score of 112 and maximum progression of 16 floors cleared, while preserving high computational efficiency with an average runtime of 158.34 seconds per step. The choice of $M=5$ strikes an ideal trade-off between behavioral diversity and execution efficiency, outperforming both smaller and larger sampling sizes. Similarly, the 1:1 reward ratio proves more effective than imbalanced configurations, better harmonizing exploration and exploitation compared to either novelty-heavy or state-value-dominated settings. Importantly, the model exhibits robust performance across a range of configurations, with multiple hyperparameter sets yielding solid results, underscoring its general insensitivity to parameter variation, though our specific configuration delivers the best balance between task advancement and operational efficiency.
@@ IV-D Model Analysis
Figure 5: Our framework executes high-reward skills even with minimal visual change between states.
Skill Evolution over Rounds. The structural evolution of the skill library and SAG across training rounds is detailed in Table IV. In Civilization V, the skill library grows dynamically through iterative augmentation and pruning, reflecting a continuous refinement process. Correspondingly, the SAG expands in complexity: node count increases from 26 to 55 and skill edges from 37 to 90, with both stabilizing after Round 2, indicating convergence toward a compact and functionally stable representation. Early rounds feature a marked rise in similarity edges, suggesting enhanced relational reasoning, followed by a slight decline by Round 3 as semantic consolidation and redundancy removal through functional merging take place. These structural improvements correlate with consistent gains in game progression and execution efficiency, accompanied by decreasing token costs, underscoring enhanced reasoning efficacy. By Round 3, convergence across all metrics confirms that repeated skill reuse and graph consolidation sustain agent adaptation and strategic exploration in open-ended settings. A parallel evolution is observed in Slay the Spire, as illustrated in Fig. 4. The structural evolution of the SAG from step 100 to step 200, with the evaluation terminating at step 210 after the agent’s character is defeated by a monster, exhibits notable expansion: nodes increase from 22 to 28, similarity edges from 190 to 312, and skill edges from 37 to 51. This growth, especially the accelerated formation of similarity edges, indicates improved recognition of state equivalences and environmental patterns. The steady rise in skill edges reflects effective knowledge transfer and progressive graph maturation, supported by denser connectivity and emergent cohesive clusters that facilitate robust decision-making.
Figure 6: Case study of successful skill invocation through state-conditioned sampling from the SAG.
Strategic Reasoning Beyond Perceptual Salience. Our framework demonstrates a consistent ability to prioritize strategic value over perceptual salience, executing high-reward skills even when visual changes between consecutive states are minimal. This capability is validated in both open-ended environments, as illustrated in Fig. 5. In Civilization V (Fig. 5(a)), the agent repeatedly selects skills such as “Advance Turn” (Total Reward: 2.038, Change_Ratio: 0.00) and “Move Worker” (Total Reward: 2.053, Change_Ratio: 0.012), which yield strategic progress despite negligible visual alterations, confirming that the reward mechanism can identify critical actions beyond superficial cues. Similarly, in Slay the Spire (Fig. 5(b)), the agent consistently employs combat skills such as “Ironclad Strike Sequence” (Total Reward: 2.114, Change_Ratio: 0.015) and “Attack and Select Strategy” (Total Reward: 2.086, Change_Ratio: 0.011), which deliver substantial rewards while producing only subtle visual differences. These selections reflect sophisticated reasoning in card sequencing, enemy targeting, and resource management—tactical decisions that cumulatively build incremental combat advantage. The ability to recognize and execute such strategically valuable actions, even in the absence of pronounced visual feedback, underscores the framework’s capacity to model complex game dynamics and pursue long-term success through sequences of subtle, high-impact decisions.
Figure 7: Case study of successful skill invocation by general VLM-guided trial and error. After five unsuccessful attempts with SAG-sampled skills, SAG-Agent switches to trial-and-error mode.
IV-E Case Study
Fig. 6 presents a case study demonstrating successful skill reuse through the SAG of SAG-Agent. When encountering a new state and corresponding node $v_{3}$, the proposed SAG-Agent first retrieves its Neighborhood of Experience $\mathcal{N}_{\mathcal{E}}(v_{3})$ and identifies high-quality candidate skills $\mathbb{S}{HQ}(v_{3})$. Skills are then sampled probabilistically according to their skill edge weights in the SAG. While the initially sampled skill “Use Defend Card in Battle” fails due to grounding issues, the subsequently sampled skill “Quick Attack Sequence” executes successfully through the simulated mouse operations “Clicked at (673, 926)” and “Clicked at (523, 951)”, achieving a total reward of 2.480 and transitioning to next state node $v_{6}$. Following successful execution, SAG-Agent updates both the fitness value and edge weight of this skill in Memory, thereby reinforcing this validated strategy for future reuse. The case study in Fig. 6 illustrates how SAG-Agent addresses the key limitations of traditional LLM-based agents in API-free environments. First, by organizing pixel-level interactions into SAG, the agent overcomes inefficient exploration by generalizing across visually distinct but functionally similar states through Neighborhood of Experience, avoiding myopic decisions. Second, the edge-weight based sampling demonstrates strategic reuse of historical knowledge rather than relying on trial-and-error. Third, the successful execution and subsequent memory update illustrate our hybrid reward mechanism: the substantial environmental reward combined with fitness updates reinforces high-value pathways while maintaining exploration flexibility. This showcases SAG-Agent’s ability to decouple strategic planning from pure discovery, effectively addressing both skill acquisition and long-horizon reasoning challenges in API-free environments.
Fig. 7 demonstrates SAG-Agent’s two-stage skill invocation strategy, which prioritizes structured knowledge from the SAG before resorting to VLM-guided search. Initially facing state node $v_{6}$, the SAG-Agent attempts SAG-based skill sampling but encounters five consecutive failures due to grounding issues or insufficient rewards. This triggers a fallback to VLM-guided skill retrieval from Procedural Memory. The process first queries the Action Clusters Table to obtain candidate skill clusters $\mathbb{S}_{C}(v_{6})=\{1,2,3,5,6,7,8,9,12\}$, from which the VLM identifies the most contextually relevant cluster based on visual observation. To navigate the stochastic environment, the SAG-Agent employs a UCT-inspired strategy that balances exploitation of high-value skills $\{\sigma_{2},\sigma_{3},\sigma_{9},\sigma_{11}\}$ with exploration of less-tried options. The resulting utility scores are converted into a probability distribution via temperature-scaled softmax, enabling stochastic sampling. The first sampled skill “Attack and Select Strategy” (Probs.=9.99e01Probs.=9.99e-01) executes successfully through operations “Clicked at (687, 933)” and “Clicked at (1134, 692)”, yielding a 2.117 reward. This successful execution triggers updates to the skill’s fitness and edge weight in memory, demonstrating SAG-Agent’s capacity for continuous knowledge refinement through both successful and failed experiences.
V Conclusion
@@
In this paper, we propose SAG-Agent to organize pixel-level GUI interactions into a knowledge graph of states and actions, enabling agents to overcome short-sightedness in API-free, open-ended environments. By linking functionally similar yet visually distinct states, the agent generalizes past experiences into coherent long-term strategies. Experiments in Civilization V and Slay the Spire demonstrate that SAG-Agent achieves significant gains in exploration efficiency and strategic depth over state-of-the-art baselines. These results highlight the potential of structuring experience into a graph-based memory, advancing API-free agents toward scalable and general-purpose autonomy.
@@
Limitations and Future Work. A key limitation of SAG-Agent is that its knowledge remains tied to specific environments, lacking the full abstraction and transferability that characterize human reasoning. While the framework structures experience into a graph-based memory, its ability to generalize knowledge across unseen tasks and diverse contexts is still constrained. Future work will focus on advancing this direction: enabling higher-level knowledge induction and reasoning that can span heterogeneous environments and ultimately testing such capabilities in more complex, real-world scenarios. This progression would move API-free agents closer to human-like adaptability and robust general intelligence.